OpenStudy (anonymous):

Determine the number of possible triangles, ABC, that can be formed given C = 85°, a = 10, and c = 13. 1 2

OpenStudy (anonymous):

@mathmath333 @triciaal

OpenStudy (anonymous):

is it one?

OpenStudy (mathmath333):

first find the value of \(\angle A\)

OpenStudy (mathmath333):

\(\large\tt \begin{align} \color{black}{\dfrac{\sin A}{a}=\dfrac{\sin C}{c}\hspace{.33em}\\~\\}\end{align}\)

OpenStudy (mathmath333):

does that makes sense ?

OpenStudy (anonymous):

yep thnx

OpenStudy (mathmath333):

so how many trinagle are possible ?